OpenStudy (openstudyoperator):

Mathew raced his bicycle 72 miles in 6 hours. What is the unit rate in miles per hour? Question 9 options: A) 9 miles per hour B) 12 miles per hour C) 6 miles per hour D) 20 miles per hour

OpenStudy (openstudyoperator):

@akiaki28

OpenStudy (steve816):

You would have to do \[\huge \frac{ 72~miles }{ 6~hours }=\frac{ 12~miles }{ 1~hour }\]

OpenStudy (akiaki28):

rate = miles/hour so 72/6 = 12 miles/hour

OpenStudy (openstudyoperator):

Ohh, thanks!

OpenStudy (steve816):

No problem! Welcome to openstudy :)